OpenStudy (anonymous):

NEED HELP SOS Find an integer, x, such that 5, 10, and x represent the lengths of the sides of an obtuse triangle. A. 4 B. 5 C. 6 D. 11

OpenStudy (welshfella):

In an obtuse triangle c^2 > a^2 + b^2

OpenStudy (welshfella):

so if we take c = 10 what inequality can we write?

OpenStudy (welshfella):

10^2 > 5^2 + b^2 can you solve this for b^2? then for b?

OpenStudy (anonymous):

100 + 25

OpenStudy (welshfella):

not +

OpenStudy (anonymous):

100 > 25 + 36 for 6

OpenStudy (anonymous):

i thought the asnwer would be 6

OpenStudy (welshfella):

b^2 < 100 - 25 b^2 < 75

OpenStudy (welshfella):

b < sqrt75 b < 8.66

OpenStudy (anonymous):

im really confused

OpenStudy (welshfella):

b cannot be 4 or 5 because that would not make a triangle and 11 > 8.66 therefore the answer is 6

OpenStudy (anonymous):

8 isnt an option

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

i have one more question

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

how do i find an acute triangle

OpenStudy (anonymous):

Using Pythagorean inequalities, determine which set of the given three sides produces an acute triangle.

OpenStudy (welshfella):

for an acute angle , if c is the longest side, c^2 < a^2 + b^2

OpenStudy (anonymous):

3, 4, 5 6,12,20 4,6,10, 5,12,18 ALL NUMBERS TO SQUARE ROOT

OpenStudy (welshfella):

- and dont forget for all triangles the sum of any 2 sides must exceed the length of the other side

OpenStudy (anonymous):

what

OpenStudy (anonymous):

i dont know how to do it cause the numbers are squared

OpenStudy (welshfella):

what exactly is the second question?

OpenStudy (anonymous):

Using Pythagorean inequalities, determine which set of the given three sides produces an acute triangle.

OpenStudy (anonymous):

6 minutes till time runs out on this question:/

OpenStudy (welshfella):

well the first one is a right angled triangle because 5^2 = 3^2 + 4^2

OpenStudy (anonymous):

what about C

OpenStudy (anonymous):

rembember all the numbers are squared

OpenStudy (welshfella):

the 2nd is not a triangle? can you see why? neither is the third why?

OpenStudy (anonymous):

4, 6,10 4x4 = 16 6x6=36

OpenStudy (welshfella):

and 10^2 = 100 but 4+6 = 10 so its not a triangle..

OpenStudy (anonymous):

oh

OpenStudy (welshfella):

you cant draw a triangle with sides this length

OpenStudy (welshfella):